OpenStudy (landshark5):

{Please help} How do you find the factors of (x-3)(x-8)?

jimthompson5910 (jim_thompson5910):

(x-3)(x-8) is completely factored The factors are simply the terms in the parenthesis

jimthompson5910 (jim_thompson5910):

So x-3 is one factor x-8 is the other

OpenStudy (landshark5):

How do you find the x values. My algebra teacher showed me but I forgot it.

OpenStudy (landshark5):

can you find the x value?

jimthompson5910 (jim_thompson5910):

well if that whole thing was set to 0, then you can use the zero product property to solve for x

jimthompson5910 (jim_thompson5910):

if (x-3)(x-8) = 0, then... (x-3)(x-8) = 0 x-3 = 0 or x-8 = 0 now solve each equation for x

OpenStudy (landshark5):

I solved it on my notebook here. The original equation was x^2+5-21

OpenStudy (landshark5):

wait let me try

jimthompson5910 (jim_thompson5910):

x^2 + 5x - 21 does NOT factor to (x-3)(x-8)

jimthompson5910 (jim_thompson5910):

notice how (-3)*(-8) = 24 and NOT -21

OpenStudy (landshark5):

I mean -24 not 21 sorry

jimthompson5910 (jim_thompson5910):

ok, notice how (-3)*(-8) = 24 which is positive, but we want negative 24

OpenStudy (landshark5):

ohh ok i get it now Thank you very much!

jimthompson5910 (jim_thompson5910):

so what does x^2 + 5x - 24 factor to?

OpenStudy (landshark5):

have a good day sir.

OpenStudy (landshark5):

well i got (x-3)(x+8) and the out come i got was x=3 or x=-8

OpenStudy (landshark5):

I believe that is how my teacher told me how to do it.